Benoit Company produces three products—A, B, and C. Data concerning the three products follow (per unit):

Product

A B C
Selling price $

85.00

$ 60.00 $ 82.00
Variable expenses:
Direct materials 27.20 15.00 8.00
Other variable expenses 27.20 33.00 53.50
Total variable expenses 54.40 48.00 61.50
Contribution margin $ 30.60 $ 12.00 $ 20.50
Contribution margin ratio 36 % 20 % 25 %

The company estimates that it can sell 850 units of each product per month. The same raw material is used in each product. The material costs $4 per pound with a maximum of 5,100 pounds available each month.

Required:

1. Calculate the contribution margin per pound of the constraining resource for each product

Product A Product B Product C
Contribution margin per pound of the constraining resource

2. Which orders would you advise the company to accept first, those for A, B, or C? Which orders second? Third?

3. What is the maximum contribution margin that the company can earn per month if it makes optimal use of its 5,100 pounds of materials?
Answer #1

1) Contributionn margin per pound

Product A Product B Product C
Contribution margin per unit 30.60 12 20.50
Pound per unit 6.8 3.75 2
Contribution margin per pound 4.5 3.2 10.25
Rank 2 3 1

2. Which orders would you advise the company to accept first, those for A, B, or C? Which orders second? Third?

PRoduct Order
A Second
B Third
C First

3)

Pound Unit
C 850*2 = 1700 850
A 3400 3400/6.8 = 500
B 0 0
Total 5100

Maximum contribution margin = (850*20.50+500*30.60) = 32725